A 12ft board is to be cut into 2 pieces. One of those pieces is 3ft longer than the other piece. What are the length of the 2 pieces?
Answer by LinnW(1048)

Set x = first length , y = second length
x = y + 3
x + y = 12
substitute y + 3 for x
(y + 3) + y = 12
2y + 3 = 12
add -3 to each side
2y = 9
y = 4.5
x = 4.5 + 3 = 7.5
7.5 + 4.5 = 12 so this checks.